5-2 Written problem: Consider two waves on two side-by-side strings along the x-axis, one of form yi(x,t) - A sin(kx - cot) and the other of form y(x, t-A sin(kx-ω1+ φ), where A-10.0 cm, k 3.0007 cm-1 and 0.516 s, as sketched below (this is a snapshot in time - the waves are moving together to the right) Problem 5-2 What value of the phase constant ф will make y2-y/2 at t-1.5 s and at x - 0 while also making the y-velocities (the up and down string velocities) of the two waves be opposite in sign at these values of x and t? Note: these velocities are not the wave velocities. These two waves have the same k and the same o, so their wave velocities down the x-axis are exactly the same. Make sure your phase angle ф is between- and .
Answer #1

We have the 2 wave equations yi = Asin(kr-wt) and y_{2}=Asin(kx-omega t+phi ) .

At t=1.5 s and x=0 we get yi = Asin (-wt)--Asinut =-Asin (1.5w) and y2 = Asin (-wt + o)--Asin(wt-o)--Asin (1.5w-o) .

Since y_{2}=y_{1}/2 at this value of x and t we must have

-Asin(1.5omega -phi )=-Asin(1.5omega )/2 i.e sin( 1.5w-o) sin(1.5w)/2

On simplifying the above equation using the trigonometric relation sin(A-B)=sin A cos B-cos A sin B we obtain:

sin(1.5w)cos - cos 1.5w)sino- sin(1.5w)/2 i.e

sin(1.5w) (coso- 1/2)-cos(1.5w)sino i.e

tan(1.5w)- sinó/(coso - 1/2)- 0.97 on substituting the value of omega and determining tan(1.5w).

i.e.sinphi =0.97 cosphi -0.49 or sinphi =0.97 (sqrt{1-cos^{2}phi })-0.49 i.e (sinphi+0.49)^{2} =0.97^{2}(1-sin^{2}phi ) . This reduces to

sin^{2}phi +0.98sinphi -0.7=0. On solving this quadratic equation we obtain :

sinphi =0.48 i.e phi =sin^{-1}(0.48)=0.5 rad or 290 .

For the y velocities to be opposite in sign at these values of x and t, we must subtract an additional value of pi from the above value of phi to obtain the value of (0.5-3.14)--2.64rad as the final value of phi.